OpenStudy (anonymous):

Multiply (2sqrt(5) - 1) * (6 sqrt(5) - 3) Helpppp :/

OpenStudy (anonymous):

\[(2\sqrt5-1)(6\sqrt5-3)\] four multiplications

OpenStudy (anonymous):

\[2\sqrt5\times 6\sqrt5-3\times 2\sqrt5-6\sqrt5+3\] \[12\times 5-12\sqrt5+3\] \[63-12\sqrt5\]

OpenStudy (anonymous):

psh thanks you!!!!